$SPAD/src/input noptip.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool noptip.output
)set message test on
)set message auto off
)clear all

———-

Note that the conversions of results from DoubleFloat to Float will become unnecessary if outputGeneral is extended to apply to DoubleFloat quantities.

— * —

--S 1 of 6
outputGeneral 5
--R
--R                                                                   Type: Void
--E 1
--S 2 of 6
f := %e^x*(4*x^2 + 2*y^2 + 4*x*y + 2*y + 1);
--R
--R
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 6
start := [x=-1.0, y=1.0];
--R
--R
--R                                      Type: List(Equation(Polynomial(Float)))
--E 3
--S 4 of 6 used to work?
nagMin(f,start) :: List Equation Polynomial Float
--R
--R   There are no library operations named nagMin
--R      Use HyperDoc Browse or issue
--R                               )what op nagMin
--R      to learn if there is any operation containing " nagMin " in its
--R      name.

--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagMin with argument type(s)
--R                             Expression(Integer)
--R                      List(Equation(Polynomial(Float)))
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 4
--       [x= 0.5,y= - 1.0]
--S 5 of 6
outputGeneral()
--R
--R                                                                   Type: Void
--E 5
--S 6 of 6
output "End of tests"
--R
--R   End of tests
--R                                                                   Type: Void
--E 6
)spool
)lisp (bye)

———-

References^{[1] nothing}

5